This page will be updated regularly

Developments to date:
| 10/22/2001 | Dual filter wheels and test collimation lenses provided by University of North Carolina. |
| 10/25/2001 | Queensgate EC50WF etalon and CS-100 controller provided by Patrick Harigan of Rice University. Preliminary optical tests show that the etalon and controller provide a stable ring pattern. |
| 11/29/2001 | IR Labs ND-8 LN2 dewar for detector testing supplied by Ball Aerospace. |
| 2/14/2002 | Placed order for the Mauna Kea Filter set JHK filters |
| 4/12/2002 | Preliminary design review. A agreement was quickly reached on the purchase of the Rockwell Hawaii 1RG 1k x 1k HgCdTe focal plane array. To our knowledge, NIC-FPS is the first ground-based instrument to use this detector. |
| 5/1/2002 | Mauna Kea Filter set H and K filters arrived from Barr Associates, Inc. |
| 7/26/2002 | Critical design review. |
| 10/22/2002 | Mounted lens assemblies from Janos have been ordered. |
| 1/28/2003 | Minutes from the telecon with Apache Point Observatory. |
| 4/4/2003 | One-year progress review, presentations listed below |
| 6/20/2003 | Preliminary fit check of the main housing goes smoothly. |
| 7/3/2003 | Initial vacuum pump down of the housing yields pressures ~ mid 10^-5 torr. Helium leak check sources the bolt holes of the main housing side doors. |
| 8/1/2003 | Cold stand off is machined and plated with iridite. |
| 10/25/2003 | Optics bench checks to be perpendicular with little deflection from rotation or weight. |
| 12/20/2003 | Heaters, charcoal getters, thermal straps and thermal sensors are installed. Mounting plate to the Nasmyth port at Apache Point Observatory is machined and fit to the front housing. |
| 1/20/2004 | NIC-FPS is currently undergoing a series of thermal tests, the first of which went quite well. |
| 4/20/2004 | Etalon and filter wheels are being tested. Internal surfaces have been painted with Aeroglaze. Optics bench is having holes drilled and components will be integrated shortly. |
| 5/1//2004 | Detector has produced its first image. |
| 5/15//2004 | Thermal blankets are finished and installed. |
| 6/5/2004 | Optical alignment test is now being setup and will commence next week. |
| 7/29/2004 | Optics, detector, filter wheels and etalon are installed and being tested in cold vacuum. |
| 9/24/2004 | The instrument leaves Casa on the first trip to Apache Point Observatory for preliminary tests. Click here for some photos. |
| 10/1/2004 | NIC-FPS returns to CASA where narrowband filter sets, and science grade detector chip are installed. Click for narrowband filter information. |
| 11/17/2004 | Instrument is installed at APO. |
| 11/21/2004 | The first series of sky images with the science grade chip are taken. Click here to view images. |
| 1/25/2005 | First remote observing done. |
| 4/5/2005 | Dewer is holding around 3 x 10^-7 torr with a 20 L/s ion pump running. |
| 7/26/2005 | Summer shutdown at Apache Point Observatory. Instrument is returned to CASA - ARL for testing new detector software and electronics, filter wheel modifications and a grism / slit installation. Etalon is still out of service. |
Publications
Presentations are PDF documents
Presentations are PDF documents
Presentations are PDF documents